# On the Curvature of Regge Metrics

### v1

@@ 2.3 Derivation of the Angle Defect
Let $\xi^{T}_{p}(x):=\eta(dF^{T}\circ d\psi_{p}^{T}(\frac{\partial}{\partial s}|_{(x,s)}))$ be the $\mathfrak{so}(n)$-valued function representing rotation around $p$ at the angular speed $r_{p,e}^{T}(x)$ (measured from the side $e$ such that the rotation from $E_{p,e}^{T}$ to $E_{p,e^{\prime}}^{T}$ is in the counterclockwise direction, where $e^{\prime}$ is the other side). Then since $\omega(X)=\eta(X)$ for any vertical vector $X$, the integral can be made simpler:
@@
at $s=0$, where $T_{j}$ and $e_{j}$ are some enumeration of the codimension-0 and codimension-1 faces surrounding $p$ such that for each $j=1,\dots,{k-1}$, $T_{j}\cap T_{j+1}=e_{j+1}$, the rotation from $E_{p,e_{j}}^{T_{j}}$ to $E_{p,e_{j+1}}^{T_{j}}$ is in the counterclockwise direction for each $j=1\dots k-1$, and $k$ is the number of codimension-1 faces incident to $p$. Another way to characterize this enumeration is that $E_{p,e_{j}}^{T_{j}}$ has an inward-pointing normal vector as its last element and $E_{p,e_{j+1}}^{T_{j}}$ has an outward-pointing normal vector as its last element. If $p$ is an interior polytope, then we also require that the preceding statements hold for $j=k$, where $T_{k+1}:=T_{1}$ and $e_{k+1}=e_{1}$.
@@
For $x\in p$, the condition 4 on the compatible frames (and the fact that $F^{T_{j}}$ is continuous) implies
$f^{T_{j}}|_{\mathring{e}_{j+1}}(x)=F^{T_{j}}(1,x)=E_{p,e_{j}}^{T_{j}}\cdot\exp\left(r_{p,e_{j}}^{T_{j}}(x)w^{n-1}_{n}\right)A_{p,e_{j}}^{T_{j}}\mu_{e_{j}}^{T_{j}}.$
However, we also know that $f^{T_{j}}|_{\mathring{e}_{j+1}}=E_{p,e_{j+1}}^{T_{j}}\cdot A_{p,e_{j+1}}^{T_{j}}\mu_{e_{j+1}}^{T_{j}}$, so because
$E_{p,e_{j+1}}^{T_{j}}=E_{p,{e_{j}}}^{T_{j}}\cdot\exp\left(\theta_{p}^{T_{j}}w^{n-1}_{n}\right)$, we can use the fact that the group action is free on each fiber to derive
@@
$\gamma(s)=sr_{p,e_{1}}^{T_{1}}+\sum_{j=2}^{k}(1+s)r_{p,e_{j}}^{T_{j}}-\theta_{p}^{T_{j}}$. For interior polytopes, this can be simplified with one more application of (13) to obtain the equivalent expression
$G_{p}(s)=\mathbf{Ad}((A_{p,e_{k}}^{T_{k}}\mu_{e_{k}}^{T_{k}})^{-1})\left(\exp\left[(\gamma(s)+(1+s)r_{p,e_{1}}^{T_{1}}-\theta_{p}^{T_{1}})w^{n-1}_{n}\right]\right).$ (14)
Therefore for interior polytopes, $G_{p}(0)=I$ implies that $\gamma(0)+r_{p,e_{1}}^{T_{1}}-\theta_{p}^{T_{1}}=2\pi m$ for an integer $m$. This quantity must be continuous with respect to continuous deformations of $g$ and $f$. Thus, using the final condition for compatible frames and the fact that $r_{p,e}^{T}=0$ for continuous frames and continuous metrics, we get m=1m=-1. Therefore $\sum_{j=1}^{k}r_{p,e_{j}}^{T_{j}}=-\Theta_{p}$, where $\Theta_{p}:=2\pi-\sum_{j=1}^{k}\theta_{p}^{T_{j}}$ is the angle defect of $p$ at $x$. Lastly we note that, due to equation (14),
@@
This expression clearly doesn’t depend on the particular enumeration $T_{j}$ we picked, so we will omit the subscript k from now on and remember that although the expression involves codimension-0 and codimension-1 polytopes $T$ and $e$ of which $p$ is a face, the expression is actually independent of the particular choice. The only caveat is that the integral needs to be evaluated with the orientation on $p$ such that $E_{p}$ is positively oriented, and $e$ needs to be chosen so that $E_{p,e}^{T}$ has an inward-pointing normal vector; otherwise the signs will be wrong.
@@ 2.4 Properties of the Distributional Curvature
Suppose $f=\bigsqcup_{T\subseteq M}f^{T}$ satisfies all of the compatibility conditions in Definition 1 and $\phi\in C^{\infty}_{c}\Omega^{n-2}_{\mathbf{t}}(M;\mathfrak{so}(n))$ is a smooth compactly supported $\mathfrak{so}(n)$-valued (n2)(n-2)-form which vanishes when pulled back to $\partial M$. Then
$\left\langle\left\langle f^{*}\Omega_{\mathrm{dist}},\phi\right\rangle\right\rangle=\sum_{T\subseteq M}\int_{\mathring{T}}$ $\left\langle{f^{T}}^{*}\Omega\wedge\phi\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$+\sum_{p\subset M}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n})\wedge i_{\mathring{p}}^{*}\phi\right\rangle.$ (16)
Furthermore, the distributional curvature is tensorial, in the following sense: if $h\in C^{\infty}(M;O(n))$, then $f\cdot h$ is a compatible frame and
@@
The proof of the expression for the distributional curvature was already carried out in subsections 2.1-2.3, and it is clear that $f\cdot h$ satisfies all the compatibility conditions if $h$ is continuous and piecewise-smooth, so all that remains to be proven is the tensoriality. First, we will expand out the expression for the distributinal curvature in the frame $\hat{f}=f\cdot h$. Note that since $f=E_{e}^{T}\cdot\mu_{e}^{T}$, $\hat{\mu}_{e}^{T}=\mu_{e}^{T}h$, so the expression is not hard to evaluate:
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle(f^{T}\cdot h)^{*}\Omega\wedge\phi\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T}h)^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$\quad+\sum_{p\subset M}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T}h)^{-1})(w^{n-1}_{n})\wedge i_{\mathring{p}}^{*}\phi\right\rangle$
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle\mathrm{Ad}(h^{-1})({f^{T}}^{*}\Omega)\wedge\phi\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left\langle\mathrm{Ad}(h^{-1})(\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!]))\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$\quad+\sum_{p\subset M}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}(h^{-1})(\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n}))\wedge i_{\mathring{p}}^{*}\phi\right\rangle$
$=\sum_{T\subseteq M}\left\langle{f^{T}}^{*}\Omega\wedge\mathrm{Ad}(h)(\phi))\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge\mathrm{Ad}(h)(i_{\mathring{e}}^{*}\phi)\right\rangle$
$\quad+\sum_{p\subset M}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n})\wedge\mathrm{Ad}(h)(i_{\mathring{p}}^{*}\phi)\right\rangle$
@@
Another observation we can make is that the right-hand side of (1) is well-defined
even if $\phi$ is a $\mathfrak{gl}(n)$-valued (n2)(n-2)-form and the frame is of the form $f\cdot h$, where $f$ is compatible and $h$ is a continuous $GL(n)$-valued map. This is useful because it allows us to consider frames which are not orthogonal but are still, in a sense, compatible.
The space of test forms can be further expanded to include discontinuous forms of a particular type.
@@
Consider a compactly supported $\mathfrak{gl}(n)$-valued (n2)(n-2)-form $\phi$ which is $C^{2}$ on $\mathring{T}$ for each codimension-0 polytope $T\subseteq M$. Such a form will be called a compatible $\mathfrak{gl}(n)$-valued test form if the following conditions are satisfied:
@@
    For each $T\subseteq M$ there exists a unique Lipschitz continuous extension of ${\Phi^{T}}^{*}\phi|_{\mathring{T}}$ to $B_{T}$, which will be called $\tilde{\phi}^{T}\in C^{0}\Omega^{n-2}(B_{T};\mathfrak{gl}(n))$. In other words, ${\Phi^{T}}^{*}\phi|_{\mathring{T}}$ has uniformly bounded first derivatives in any coordinate chart. This implies that $i_{\mathring{e}}^{*}\phi|_{T}={\Phi^{T}}^{-*}i_{{\Phi^{T}}^{-1}(\mathring{e})}^{*}\tilde{\phi}^{T}$ is a well-defined continuous extension of $\phi|_{\mathring{T}}$, and it implies that if $d\Phi^{T}(v)=0$, then $v\lrcorner\tilde{\phi}^{T}=0$ (this is relevant only on the boundary of $B_{T}$).
@@
    If $e=T\cap T^{\prime}$ is a codimension-1 interface, then $i_{e}^{*}\phi|_{T}=i_{e}^{*}\phi|_{T^{\prime}}$. In other words, $i_{e}^{*}\phi]\!]=0$.
@@
    If $p\subset T$ is a codimension-2 face, for each $s\in[0,1]$ define the map $\chi_{s,p}^{T}:p\to\overline{{\Phi^{T}}^{-1}(\mathring{p})}$ by $\chi_{s,p}^{T}(x)=\psi_{p}^{T}(s,x)$. Then ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}={\chi_{0,p}^{T}}^{*}\tilde{\phi}^{T}$ for any $s\in[0,1]$. In other words, $i_{\mathring{p}}^{*}\phi$ is well-defined, continuous, and bounded on $\mathring{p}$.
@@
If $\phi$ is a compatible $\mathfrak{gl}(n)$-valued test form, then the arguments in Section 2.2-2.3 remain valid with essentially no modification. A notable example of compatible test forms are the forms $\frac{1}{2}\phi_{ij}\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$, where each $\phi_{ij}:M\to\mathfrak{gl}(n)$, $i,j=1,2,\dots,n$, is a compactly supported continuous map that is $C^{2}$ on the interior of each codimension-0 polytope and vanishes on $\partial M$. The map $\star:\Omega^{2}(M)\to\Omega^{n-2}(M)$ is, on each codimension-0 polytope $T$, the Hodge star operator associated to the metric $g_{T}$.
We will define a generalized compatible frame to be a frame $f$ such that there exists a continuous map $h:M\to GL(n)$ where $f\cdot h$ is a compatible frame. Equation (1) can then be taken as a definition when $f$ is a generalized compatible frame and $\phi$ is a compatible $\mathfrak{gl}(n)$-valued test form.
Remark 2.
We have ignored the question of whether the piecewise-smooth one-form $f^{*}\omega$ defined by ${f^{*}\omega}^{i}_{j}|_{T}=\left\langle\nabla^{T}f_{j}^{T},f_{i}^{T}\right\rangle_{T}$ on $T$ is “really” the connection form associated to the compatible frame $f$, despite the fact that $f$ is discontinuous. Here is one way to answer that question: Since the pullback of $f^{*}\theta^{i}$ to every element interface $e=T\cap T^{\prime}$ is single-valued, the distributional exterior derivative of $f^{*}\theta^{i}$ is simply its elementwise exterior derivative. Therefore the structure equations $df^{*}\theta^{i}=-f^{*}\omega^{i}_{j}\wedge f^{*}\theta^{j}$ hold in a distributional sense.

### v2

@@ 2.3 Derivation of the Angle Defect
Let $\xi^{T}_{p}(x):=\eta(dF^{T}\circ d\psi_{p}^{T}(\frac{\partial}{\partial s}|_{(x,s)}))$ be the $\mathfrak{so}(n)$-valued function representing counterclockwise rotation around $p$ at the angular speed $r_{p,e}^{T}(x)$. The fact that $\xi$ is independent of $s$ is a consequence of condition 3. Then since $\omega(X)=\eta(X)$ for any vertical vector $X$, the integral can be made simpler:
@@
at $s=0$, where $T_{j}$ and $e_{j}$ are some enumeration of the codimension-0 and codimension-1 faces surrounding $p$ such that for each $j=1,\dots,{k-1}$, $T_{j}\cap T_{j+1}=e_{j+1}$, the rotation from $E_{p,e_{j}}^{T_{j}}$ to $E_{p,e_{j+1}}^{T_{j}}$ is in the counterclockwise direction for each $j=1,\dots,k-1$, and $k$ is the number of codimension-1 faces incident to $p$. Another way to characterize this enumeration is that $E_{p,e_{j}}^{T_{j}}$ has an inward-pointing normal vector as its last element and $E_{p,e_{j+1}}^{T_{j}}$ has an outward-pointing normal vector as its last element. If $p$ is an interior polytope, then we also require that the preceding statements hold for $j=k$, where $T_{k+1}:=T_{1}$ and $e_{k+1}:=e_{1}$.
@@
For $x\in p$, condition 3 on the compatible frame (and the fact that $F^{T_{j}}$ is continuous) implies
$f^{T_{j}}|_{e_{j+1}}(x)=F^{T_{j}}(\psi_{p}^{T_{j}}(1,x))=E_{p,e_{j}}^{T_{j}}\cdot\exp\left(r_{p,e_{j}}^{T_{j}}(x)w^{n-1}_{n}\right)A_{p,e_{j}}^{T_{j}}\mu_{e_{j}}^{T_{j}}.$
However, we also know that $f^{T_{j}}|_{e_{j+1}}=E_{p,e_{j+1}}^{T_{j}}\cdot A_{p,e_{j+1}}^{T_{j}}\mu_{e_{j+1}}^{T_{j}}$, so because the rotation from $E_{p,e_{j}}^{T_{j}}$ to $E_{p,e_{j+1}}^{T_{j}}$ is in the counterclockwise direction, we have $E_{p,e_{j+1}}^{T_{j}}=E_{p,{e_{j}}}^{T_{j}}\cdot\exp\left(\theta_{p}^{T_{j}}w^{n-1}_{n}\right)$, so we can use the fact that the group action is free on each fiber of $\mathcal{F}_{O}(T_{j})$ to derive
@@
$\gamma(s):=sr_{p,e_{k}}^{T_{k}}+\sum_{j=1}^{k-1}(1+s)r_{p,e_{j}}^{T_{j}}-\theta_{p}^{T_{j}}.$
For interior polytopes, this can be simplified with one more application of (13) to obtain the equivalent expression
$G_{p}(s)=\mathbf{Ad}((A_{p,e_{k}}^{T_{k}}\mu_{e_{k}}^{T_{k}})^{-1})\left(\exp\left[(\gamma(s)+r_{p,e_{k}}^{T_{k}}-\theta_{p}^{T_{k}})w^{n-1}_{n}\right]\right).$ (14)
Therefore for interior polytopes, $G_{p}(0)=I$ implies that $\gamma(0)+r_{p,e_{k}}^{T_{k}}-\theta_{p}^{T_{k}}=2\pi m$ for an integer $m$. The integer $m$ must be continuous with respect to continuous deformations of $g$ and $f$. Thus, using the final condition for compatible frames and the fact that $r_{p,e}^{T}=0$ for continuous frames and continuous metrics, we get m=1m=-1. Therefore $\sum_{j=1}^{k}r_{p,e_{j}}^{T_{j}}=-\Theta_{p}$, where $\Theta_{p}:=2\pi-\sum_{j=1}^{k}\theta_{p}^{T_{j}}$ is the angle defect of $p$ at $x$. Lastly we note that, due to equation (14),
@@
The left-hand side of this equation clearly doesn’t depend on the particular enumeration $T_{j}$ we picked, so we will omit the subscript k from now on and remember that although the expression on the right-hand side involves particular codimension-0 and codimension-1 polytopes $T$ and $e$ of which $p$ is a face, the expression is actually independent of the particular choice. The only caveat is that the integral needs to be evaluated with the orientation on $p$ induced from $e$, and $e$ needs to be chosen so that the last entry of $E_{p,e}^{T}$ is an inward-pointing normal vector; otherwise the sign on $\Theta_{p}$ would need to flip.
@@ 2.4 Properties of the Distributional Curvature
Suppose $f=\bigsqcup_{T\subseteq M}f^{T}$ satisfies all of the compatibility conditions in Definition 2 and $\phi\in C^{\infty}_{c}\Omega^{n-2}(M;\mathfrak{so}(n))$ is a smooth compactly supported $\mathfrak{so}(n)$-valued (n2)(n-2)-form which vanishes when pulled back to $\partial M$. Then
$\left\langle\left\langle f^{*}\Omega_{\mathrm{dist}},\phi\right\rangle\right\rangle=\sum_{T\subseteq M}\int_{\mathring{T}}$ $\left\langle{f^{T}}^{*}\Omega\wedge\phi\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n})\wedge i_{\mathring{p}}^{*}\phi\right\rangle.$ (16)
Furthermore, the distributional curvature is tensorial, in the following sense: if $h$ is a continuous piecewise-smooth map $M\to O(n)$, then $f\cdot h$ is a compatible frame and
@@
The proof of the expression for the distributional curvature was already carried out in subsections 2.1-2.3, and it is clear that $f\cdot h$ satisfies all the compatibility conditions if $h$ is continuous and piecewise-smooth, so all that remains to be proven is the tensoriality. First, we will expand out the expression for the distributional curvature in the frame $\hat{f}=f\cdot h$. Note that since $f=E_{e}^{T}\cdot\mu_{e}^{T}$, $\hat{\mu}_{e}^{T}=\mu_{e}^{T}h$, so the expression is not hard to evaluate:
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle(f^{T}\cdot h)^{*}\Omega\wedge\phi\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T}h)^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$\quad+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T}h)^{-1})(w^{n-1}_{n})\wedge i_{\mathring{p}}^{*}\phi\right\rangle$
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle\mathrm{Ad}(h^{-1})({f^{T}}^{*}\Omega)\wedge\phi\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left\langle\mathrm{Ad}(h^{-1})(\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!]))\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$\quad+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}(h^{-1})(\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n}))\wedge i_{\mathring{p}}^{*}\phi\right\rangle$
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle{f^{T}}^{*}\Omega\wedge\mathrm{Ad}(h)(\phi))\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge\mathrm{Ad}(h)(i_{\mathring{e}}^{*}\phi)\right\rangle$
$\quad+\sum_{\mathring{p}\subset\mathring{M}}\int_{\mathring{p}}\Theta_{p}\left\langle\mathrm{Ad}((A_{p,e}^{T}\mu_{e}^{T})^{-1})(w^{n-1}_{n})\wedge\mathrm{Ad}(h)(i_{\mathring{p}}^{*}\phi)\right\rangle$
@@
Another observation we can make is that the right-hand side of (4) evaluates to zero if $\phi$ is instead $\mathrm{sym}_{n\times n}(\mathbb{R})$-valued, since all of the integrals are against $\mathfrak{so}(n)$-valued forms. This supports the intuitive notion that $f^{*}\Omega_{\mathrm{dist}}$ is $\mathfrak{so}(n)$-valued”, as $\mathfrak{so}(n)$ and $\mathrm{sym}_{n\times n}(\mathbb{R})$ are orthogonal under the bilinear product we are using.
The space of test forms can also be expanded to include discontinuous forms of a particular type.
@@
Consider a compactly supported $\mathfrak{so}(n)$-valued (n2)(n-2)-form $\phi$ which is $C^{2}$ on $\mathring{T}$ for each codimension-0 polytope $T\subseteq M$. Such a form will be called a compatible $\mathfrak{so}(n)$-valued test form if the following conditions are satisfied:
@@
    For each $T\subseteq M$ there exists a unique Lipschitz continuous extension of ${\Phi^{T}}^{*}\phi|_{\mathring{T}}$ to $B_{T}$, which will be called $\tilde{\phi}^{T}\in C^{0}\Omega^{n-2}(B_{T};\mathfrak{so}(n))$. In other words, ${\Phi^{T}}^{*}\phi|_{\mathring{T}}$ has uniformly bounded first derivatives in any coordinate chart. This implies $i_{\mathring{e}}^{*}\phi|_{T}:={\Phi^{T}}^{-*}i_{{\Phi^{T}}^{-1}(\mathring{e})}^{*}\tilde{\phi}^{T}$ is a well-defined continuous extension of $\phi|_{\mathring{T}}$. It also implies that if $d\Phi^{T}(v)=0$, then $v\lrcorner\tilde{\phi}^{T}=0$ (this is relevant only on the boundary of $B_{T}$).
@@
    If $e=T\cap T^{\prime}$ is a codimension-1 interface, then $i_{\mathring{e}}^{*}\phi|_{T}=i_{\mathring{e}}^{*}\phi|_{T^{\prime}}$. In other words, $i_{\mathring{e}}^{*}\phi]\!]=0$.
@@
    The form ${\psi_{p}^{T}}^{*}\tilde{\phi}^{T}$ is basic for the fibration $q:[0,1]\times p\to p$. More precisely, if $\chi_{s,p}^{T}(x):=\psi_{p}^{T}(s,x)$, Then ${\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}={\chi_{0,p}^{T}}^{*}\tilde{\phi}^{T}$ for any $s\in[0,1]$. In other words, $i_{\mathring{p}}^{*}\phi|_{T}$ is well-defined and continuous on $p$, although $\phi$ itself may be discontinuous at $p$.
@@
If $\phi$ is a compatible $\mathfrak{so}(n)$-valued test form, then the proofs in Sections 2.2-2.3 remain valid with the small modification of using $\tilde{\phi}^{T}$ in place of ${\Phi^{T}}^{*}\phi$. Note that items (2) and (3) together imply that $i_{\mathring{p}}^{*}\phi={\chi_{s,p}^{T}}^{*}\tilde{\phi}^{T}$ is well-defined independently of $s$ and $T$. A notable example of compatible test forms are the forms $\phi_{ij}\otimes\star f^{*}(\theta^{i}\wedge\theta^{j})$, where each $\phi_{ij}:M\to\mathfrak{so}(n)$, $i,j=1,2,\dots,n$, is a compactly supported continuous map that is $C^{2}$ on the interior of each codimension-0 polytope and vanishes on $\partial M$. The map $\star:\Omega^{2}(M)\to\Omega^{n-2}(M)$ is, on each codimension-0 polytope $T$, the Hodge star operator associated to the metric $g^{T}$.